OpenStudy (anonymous):

Choose the system of equations which matches the following graph: a line includes points 0 comma 0 and negative 1 comma 2. A line includes points 2 comma 3 and 1 comma negative 2 5x - y = 7 2x - y = 0 5x - y = 7 2x + y = 0 5x + y = -7 2x + y = 0 5x + y = -7 2x - y = 0

OpenStudy (anonymous):

Line 1 = (0,0) (-1,2) Line 2 = (2,3) (1,-2) line 1 gradient: \[m = \frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\] \[m = \frac{ 2 - 0 }{ -1-0 } = \frac{ 2 }{ -1 }= -2\] \[\frac{ y - y _{1} }{ y _{2}-y _{1} }=\frac{ x - x _{1} }{ x _{2}- x _{1} }\] \[\frac{ y -0 }{2- 0 } = \frac{ x - 0 }{ - 1 - 0 } \] \[\frac{ y }{ 2 } = \frac{ x }{ -1 } \] \[-y = 2x\]

OpenStudy (anonymous):

Line 2 gradient: \[m = \frac{ y _{2}-y _{1} }{ x _{2} - x _{1} }\] \[m = \frac{ -2 - 3 }{ 1-2 } = \frac{ -5 }{ -1 } = 5\] \[y - y _{1} = m(x - x _{1})\] \[y - 3 = 5(x-2)\] \[y = 5x - 7\]

OpenStudy (anonymous):

5x - y = 7 2x - y = 0